# PhyBlock: A Progressive Benchmark for Physical Understanding and Planning via 3D Block Assembly

### v1

@@ B.2 Two Evaluation Settings for 3D Block Assembly Step Planning
A. Pose-Constrained Evaluation (with Euler Pose Consideration).
Under this strict setting, the model must produce an assembly sequence that exactly matches the ground truth in 3D Euler space (i.e., orientation in XYZ rotational angles), as detailed in Algorithm 1. Each predicted step is considered correct only if the selected block’s shape, color, and Euler pose all match the corresponding ground-truth attributes. This paradigm emphasizes the model’s geometric reasoning ability, particularly its precision in understanding 3D rotational configurations. As shown in Fig. 12, even a minor pose error in the yellow cuboid leads to structural failure in subsequent steps, highlighting the setting’s sensitivity to long-range planning inconsistencies. This strict constraint enables a fine-grained evaluation of the model’s robustness in complex spatial reasoning tasks.
B. Topology-Oriented Evaluation (without Euler Pose Consideration).
To decouple pose sensitivity from structural planning performance, this relaxed setting ignores Euler pose differences. A predicted step is deemed correct if the shape and color of the selected block match those of the ground truth, regardless of its pose. This paradigm focuses on the model’s structural planning and task decomposition capabilities, effectively mitigating the impact of local pose inaccuracies on global evaluation. As shown in Fig. 12, although the yellow cuboid has an incorrect pose, it still functions as a supporting structure, allowing the assembly to proceed correctly. This setting is particularly well-suited for evaluating the model’s high-level reasoning and planning competence.
@@ B.3 Four Error Type Classifications for 3D Block Assembly Step Analysis
To better understand the limitations of multimodal large models in multi-step reasoning tasks for 3D block assembly, we introduce a structured error analysis framework that categorizes step-level mistakes into four types: Euler Error, Dependency Not Met, Shape Not in GT, and Overflow.
Euler Error.
This error occurs when the model correctly identifies the block type but predicts an incorrect orientation, specifically in terms of Euler angles. Such rotation errors lead to misaligned placements despite the correct choice of block.
@@
Figure 13 presents illustrative examples for each of the defined error types. Specifically, in Assembly Trace (1), Step 3 involves a redundant placement of the red arch block, which is categorized as an Overflow Error. In Assembly Trace (2), Step 3 incorrectly selects an unnecessary orange cylinder that does not belong to the target block set, constituting a Shape Not in GT error. In Step 4, although the yellow cuboid is the correct block, its predicted orientation does not match the ground-truth pose, thus falling under the Euler Error category. Moreover, the failure to properly place the yellow cuboid in the second layer results in a cascade of Dependency Not Meet errors, as blocks in the third, fourth, and fifth layers rely on it for support and cannot be placed successfully.
@@ Appendix D More Results and Analysis
D.1 More Experiment Findings

### v2

@@ B.2 Two Evaluation Settings for 3D Block Assembly Step Planning
A. Pose-Constrained Evaluation (with orientation Consideration).
Under this strict setting, the model must produce an assembly sequence that exactly matches the ground truth orientation, as detailed in Algorithm 1. Each predicted step is considered correct only if the selected block’s shape, color, and orientation all match the corresponding ground-truth attributes. This paradigm emphasizes the model’s geometric reasoning ability, particularly its precision in understanding 3D rotational configurations. As shown in Fig. 12, even a minor pose error in the yellow cuboid leads to structural failure in subsequent steps, highlighting the setting’s sensitivity to long-range planning inconsistencies. This strict constraint enables a fine-grained evaluation of the model’s robustness in complex spatial reasoning tasks.
B. Topology-Oriented Evaluation (without orientation Consideration).
To decouple pose sensitivity from structural planning performance, this relaxed setting ignores the orientation differences. A predicted step is deemed correct if the shape and color of the selected block match those of the ground truth, regardless of its pose. This paradigm focuses on the model’s structural planning and task decomposition capabilities, effectively mitigating the impact of local pose inaccuracies on global evaluation. As shown in Fig. 12, although the yellow cuboid has an incorrect pose, it still functions as a supporting structure, allowing the assembly to proceed correctly. This setting is particularly well-suited for evaluating the model’s high-level reasoning and planning competence.
@@ B.3 Four Error Type Classifications for 3D Block Assembly Step Analysis
To better understand the limitations of multimodal large models in multi-step reasoning tasks for 3D block assembly, we introduce a structured error analysis framework that categorizes step-level mistakes into four types: Incorrect orientation chosen, Dependency Not Met, Shape Not in GT, and Overflow.
Incorrect Orientation Chosen.
This error occurs when the model correctly identifies the block type but predicts an incorrect orientation. Such rotation errors lead to misaligned placements despite the correct choice of block.
@@
Figure 13 presents illustrative examples for each of the defined error types. Specifically, in Assembly Trace (1), Step 3 involves a redundant placement of the red arch block, which is categorized as an Overflow Error. In Assembly Trace (2), Step 3 incorrectly selects an unnecessary orange cylinder that does not belong to the target block set, constituting a Shape Not in GT error. In Step 4, although the yellow cuboid is the correct block, its predicted orientation does not match the ground-truth pose, thus falling under the Incorrect Orientation Chosen category. Moreover, the failure to properly place the yellow cuboid in the second layer results in a cascade of Dependency Not Meet errors, as blocks in the third, fourth, and fifth layers rely on it for support and cannot be placed successfully.
@@ Appendix D More Results and Analysis
Repeated evaluations with five random seeds were conducted to ensure robust and reliable model performance. This approach guarantees that the reported outcomes are not influenced by random fluctuations and provides a more consistent measure of the models’ performance. For each model, the mean and standard deviation (±) of the F1 score were computed across various difficulty levels, from Level 1 to Level 4, as well as the overall performance. The evaluation includes both closed-source models, such as GPT-4.1, and open-source models, such as Qwen2.5-VL-7B-Instruct and Cosmos-Reason1-7B. The performance of these models is summarized in Table 5.
Table 5: F1 Scores for Model Evaluation across Difficulty Levels.
[…]
Level 1 Level 2 Level 3 Level 4
GPT-4.1 94.42 ± 2.3 46.26 ± 2.1 39.13 ± 2.4 36.07 ± 1.3 39.73 ± 1.1
Qwen2.5-VL-7B-Instruct 44.13 ± 1.3 24.12 ± 1.1 21.12 ± 2.1 16.71 ± 1.7 20.40 ± 1.2
Cosmos-Reason1-7B 43.45 ± 1.8 29.42 ± 2.1 21.53 ± 1.7 20.42 ± 1.5 23.27 ± 1.2
As shown in Table 5, GPT-4.1 consistently outperforms the open-source models across all levels, with notably low standard deviations (±2.3 to ±1.1), indicating that its performance is stable and reproducible. In contrast, the open-source models exhibit larger variations in their results, suggesting less consistency across different runs. These low error margins for GPT-4.1 confirm the reliability of its performance and demonstrate that the observed differences are not due to random fluctuations. This reinforces the robustness of our evaluation framework and underscores the superior reasoning ability of GPT-4.1, particularly when compared to the open-source models.
The comparable F1 scores observed between Level 3 and Level 4 indicate a strong transfer of reasoning patterns across structural complexities. Upon deeper inspection, our extendable data creation logic involves constructing (partially) Level 4 structures by horizontally merging two different Level 3 structures, while maintaining the same vertical depth. This design allows the reasoning patterns learned for Level 3 to transfer effectively to Level 4, enabling the model to reuse similar inference chains. However, the merging process also introduces spatial reasoning challenges, such as partial visual occlusions and extended reasoning chains, which may compound earlier errors and cause minor inconsistencies. These subtle factors likely explain the observed performance plateau and slight variations across Levels 3 and 4.
@@ E.4 Pose Estimation as a Bottleneck for 3D Assembly Tasks
E.5 Scope of Evaluation and Sim-to-Real Considerations
PhyBlock is intentionally designed to evaluate two core competencies of Vision-Language Models: (i) physical and spatial perception, and (ii) high-level assembly planning. Our evaluation setup does not currently include real-robot experiments. Incorporating real-world manipulation would introduce additional challenges, such as grasp synthesis, trajectory optimization, calibration, and hardware reliability. These factors could obscure the benchmark’s diagnostic clarity, making it difficult to determine whether performance limitations stem from reasoning or from actuation. By isolating perception and planning within a physics-accurate simulator, PhyBlock provides precise and reproducible measurements of a model’s reasoning abilities. Nonetheless, we recognize that bridging the sim-to-real gap remains an important future direction. Notably, the action sequences generated in Genesis are already compatible with standard robotic manipulation stacks (e.g., MoveIt and Cartesian impedance controllers), facilitating potential transfer to real-world robotic systems.